OpenStudy (anonymous):

Solving Radical Equations √x+2(end of sqrt)+4=x

OpenStudy (anonymous):

The first thing that comes to mind is to subtract 4 from each side so that the radical is by itself on the left side. So do this first.

OpenStudy (anonymous):

\[\sqrt{x+2} + 4 = x\] is it like this

OpenStudy (anonymous):

Next we should square each side to get rid of the radical. You are going to have to do (x-4)(x-4) on the right

OpenStudy (anonymous):

Now that teh radical is gone, we can group like terms, set it equal to zero, and either factor of use the quadratic to solve for x

OpenStudy (anonymous):

or**

OpenStudy (anonymous):

\[\sqrt{x+2} = x-4\] Square Both Sides \[x+2 = (x-4)^2\]

OpenStudy (anonymous):

\[x+ 2 = x^2 -8x + 16\] \[x^2 - 9x +14 = 0\] Now Find X usind Quadratic Formula